Iwas Realty is a real estate company selling homes in the Vancouver, British Columbia.
The mean number of homes sold by the company per day is 2.
a. What is the probability that the company sells exactly 3 homes tomorrow?
b. Find the probability of selling more than 2 homes on a particular day?

Found 2 solutions by Boreal, ewatrrr:
Answer by Boreal(15235)

this is likely a Poisson distribution with parameter proportion to time, discrete variable, and theoretically infinite numbers.
p(3)=e^(-2)*2^3/3!=0.1804
-
this is 1- probability (0,1,2)
prob of 0 is e^(-2)=0.1353
p(1) is e^(-2)*2^1/1=0.2707
p(2)=e^(-2)*2^2/2!=0.2707
That probability sum is 0.6767
the answer of p(>2) is 0.3233
also on calculator 1-poissoncdf(2,2) for the left side of the distribution.

Answer by ewatrrr(24785)

 
Hi,

The mean number of homes sold by the company per day is 2

Poisson Distribution:  λ = 2  
Using TI Calculator: 

a. What is the probability that the company sells exactly 3 homes tomorrow?
   P(3) = poissonpdf(λ,x) = poissonpdf(2,3) = .1804

b. Find the probability of selling more than 2 homes on a particular day?
      1 - P(x ≤ 2 ) = 1 - poissoncdf(2,2)=  1 - .6767 =  .3233

Personally - use a Casio fx-115 ES plus 
Important You are comfortable using Your calculator

Wish You the Best in your Studies.